# Enhanced Conformal BMS3 Symmetries

### v1

@@ 1 Introduction
The conformal BMS3 algebra seems to be very rigid, since the central extensions and the coefficients of the nonlinear terms become entirely determined by the central charge of the Virasoro subalgebra. Indeed, the conditions obtained from the Jacobi identity turn out to be very stringent, which suggests that the algebra is unique. In this sense, since the conformal BMS3 algebra looks undeformable, one may wonder whether it might be enhanced in some appropriate way. As an strategy to achieve this task we propose exploring the asymptotic structure of a suitable extension of conformal gravity in 3D vanNieuwenhuizen:1985cx ; Horne:1988jf . A nice and simple theory enjoying the sought features was proposed long ago by Pope and Townsend Pope:1989vj with the aim of further enlarging it in order to describe an infinite tower of conformal higher spin fields in 3D, along the lines of Fradkin:1987ks . More recently, conformal gravity was shown to admit a different extension that accommodates a large class of theories with a finite number of conformal higher spin fields Grigoriev:2019xmp . [^1] Additional interesting results concerning conformal higher spin fields can be found in e.g., Henneaux:2015cda ; Lovrekovic:2024yoo ; Diaz:2024iuz .
@@ 2 Asymptotic structure of extended conformal gravity in 3D
Following Pope and Townsend Pope:1989vj the three-dimensional conformal algebra $so(3,2)$, spanned by $\left\{J_{a},P_{a},K_{a},D\right\}$ with $a=0,1,2$, is enhanced to that of $so(4,2)$ by enlarging the set of generators to include $\left\{U_{a},U,V\right\}$. This is also isomorphic to algebra of depth-two conformal gravity in Grigoriev et al. Grigoriev:2019xmp .
@@ 2.1 Boundary conditions and enhanced conformal BMS3 algebra
The generators of the asymptotic symmetries can then be straightforwardly obtained following different approaches, as in Regge:1974zd ; Barnich:2001jy , and they are given by
@@ 3 Ending remarks
It is worth noting that the enhanced conformal BMS3 algebra (13) can also be seen as an infinite-dimensional nonlinear extension of the AdS5 algebra with nontrivial central charges. Thus, the obstruction to include non trivial central extensions for semisimple algebras, supported by a classical theorem of algebraic cohomology (see e.g. Fuks ), can be circumvented due to the nonlinearity of the algebra.

### v2

@@ 1 Introduction
The conformal BMS3 algebra seems to be very rigid, since the central extensions and the coefficients of the nonlinear terms become entirely determined by the central charge of the Virasoro subalgebra. Indeed, the conditions obtained from the Jacobi identity turn out to be very stringent, which suggests that the algebra is unique. In this sense, since the conformal BMS3 algebra looks undeformable, one may wonder whether it might be enhanced in some appropriate way. As a strategy to achieve this task we propose exploring the asymptotic structure of a suitable extension of conformal gravity in 3D vanNieuwenhuizen:1985cx ; Horne:1988jf . A nice and simple theory enjoying the sought features was proposed long ago by Pope and Townsend Pope:1989vj with the aim of further enlarging it in order to describe an infinite tower of conformal higher spin fields in 3D, along the lines of Fradkin:1987ks . More recently, conformal gravity was shown to admit a different extension that accommodates a large class of theories with a finite number of conformal higher spin fields Grigoriev:2019xmp . [^1] Additional interesting results concerning conformal higher spin fields in 3D can be found in e.g., Nilsson:2013tva ; Henneaux:2015cda ; Basile:2017mqc ; Kuzenko:2020opc ; Diaz:2024iuz ; Lovrekovic:2024yoo ; Grigoriev:2020lzu .
@@ 2 Asymptotic structure of extended conformal gravity in 3D
Following Pope and Townsend Pope:1989vj the three-dimensional conformal algebra $so(3,2)$, spanned by $\left\{J_{a},P_{a},K_{a},D\right\}$ with $a=0,1,2$, is enhanced to that of $so(4,2)$ by enlarging the set of generators to include $\left\{U_{a},U,V\right\}$. This is also isomorphic to the algebra of depth-two conformal gravity in Grigoriev et al. Grigoriev:2019xmp .
@@ 2.1 Boundary conditions and enhanced conformal BMS3 algebra
The generators of the asymptotic symmetries can then be straightforwardly obtained following different approaches, as in Regge:1974zd ; Barnich:2001jy (see also e.g., Balachandran:1991dw ; Banados:1994tn ; Carlip:2005zn ), and they are given by
@@ 3 Ending remarks
It is worth noting that the enhanced conformal BMS3 algebra (13) can also be seen as an infinite-dimensional nonlinear extension of the AdS5 algebra with nontrivial central charges [^3] An infinite-dimensional linear extension of AdS5 has been proposed in Compere:2019bua ; Fiorucci:2020xto .. Thus, the obstruction to include non trivial central extensions for semisimple algebras, supported by a classical theorem of algebraic cohomology (see e.g. Fuks ), can be circumvented due to the nonlinearity of the algebra.